Jun 15-16

Summary

Before the sunset heavy clouds over the dome.
At the sun set clouds at the horizon.
The sky is clear after the LUCI switch with low wind and seeing 0.7-1.0.
The LUCI detector has noise pattern related to problem with the readout (the reason is stil unknown). The pattern is even variable: sometimes there are just a cuople of stripes in top right quadrants, and some other all the quadrants are strongly affected.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard hz44 red grating -  
Standard hz44 dual grating -  
Bellazzini DDT HVC274.68+74 red grating 1h mostly within seeing specs
Beltran_01 A6 H2/Ks 50m/7.5m within seeing specs
Beltran_01 A1 H2/Ks 50m/7.5m within seeing specs
Telluric FS-148 H2/Ks    

Jun 16-17

Summary

Before the sunset heavy clouds .
After twilight clear sky, with low wind and humidity.
At the beginning of scientific operations the average seiing on the GCS is 1.0.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard hz44 red grating - airmass 1.02
Standard hz44 dual grating -  
Bellazzini DDT HVC274.68+74 red grating 1h  
Beltran_01 A2_G35.03+0.35 H2, Ks 50m, 7.5m within seeing specs. Target Completed
Magrini_24 7C1756_Mask1 210_zJHK+H 1h20m mostly out of seeing specs+clouds passing. 1h to be repeated

Jun 17-18

Summary

At the beginning of the night we are closed due to heavy clouds with chance of rain and strong wind gusts.
When we open the dome the sky is only partly cloudy, the seeing is either bad and gets better only at the end of the night when the humidity increases.
We had condensation on LUCI during all the night.
Guide prove error.
Flexure compensation error
MOS Acquisition Gui got stuck.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Beltran_01 A4 H2 32m out of specs, humidity
Magrini_24 Mask1 210_zJHK, H 50min out of specs, no guiding

Jun 18-19

Summary

At the beginning of the nigh we have clear sky all over the orison, now wind and low humidity.
The average seeing at the beginning of scientific operations is 1.0 on the GCS.
Thin cirrus passing during the 1st LUCI integration.
With LUCI we got a flexure compensation error and the WFS was not working for few minutes thus an offset failed during acquisition.
AGw got stuck imaging acquisition.
The readout problem persist and get worse at the end of the night.
We also got condensation on the camera

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_46_DDT HVC274.68+74 red gradint 1h within the specs PROGRAM COMPLETED
Standard HIP82884 210_zJHK, J-band   thin cirrus, to be repeated
Magrini_24 7C1756 Mask1 210_zJHK, J-band 150min LUCI bias+pickup noise
Cresci_8 campoB Ks 26m LUCI bias+pickup noise. Bad WFS corrections, too close to the Moon

Jun 19-20

Summary

At the enclosure opening we have a mostly clear sky (few clouds remaining on the West horizon), no wind and 40% humidity
When scientific operations start we have clear sky with seeing 0.8.
Some difficulties to collimate with LBC for the first target, too low and with still too high illumination from the twilight. We got and error from the blue filter wheel, it happened twice. It was fixed by running again the ob.
Major issues from MODS: stuck of the readout. Reset of the software. Restart of the software. Server not responding --> Restart of the computer. All our, SA and TO efforts failed. Art the end of the night the problem is still unknown. An issue track was send and both the IS (O. Khun) and the PI (R. Pogge) have been notified.
Since we can not run anymore LBC observation due to the Moon, we switch back to LUCI.
The readout noise in LUCI was affecting only the upper right quadrant while acquiring the first target. The effect increased during the acquisition on the last one.
While observing with LUCI the seeing measured on Ks-H2 images is 0.45-0.6
We got the probe error with LUCI, few images are without WFS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_6Z HVC26.01 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6W HVC28.07 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec PROGRAM COMPLETED
Beltran1_A7 G31.41+0.31 H2, Ks 50min, 8min COMPLETED
Beltran1_A1 G35.20-0.74N H2, Ks 10min, 1.6min partially repeated to update the data from the 1st night

Jun 20-21

Summary

At enclosure opening the sky is clear with no wind, and 35% humidity
At the beginning of scientific operations the seeing is 1.2 on the GCS and thin cirrus on the orison
The IMCS timed our several times. We calla the SA that increase the laser power, this fails again several times.
Diffuculties to collimate: the procedure converge but there is strong coma aberration, we have to sent +2000 Z8 to the right mirror During LBC observations cirrus are passing

Data Summary

Project_RUN Objects Filter ExpTime Notes
Gilli4_A7 FIRST-J1427 r-Sloan, i-Sloan 30min, 30min out of seeing specs
Gilli4_A7 FIRST-J1427 r-Sloan, z-Sloan 30min, 30min out of seeing specs
Israel28_A2 CATS J184441 dual grating 1h 30min Target COMPLETED
Israel28_A2 Koppf27 dual grating 15 s standard required by the PI
Israel28_A5 CATS J215447 dual grating 30min mostly within seeing specs

Jun 21-22

Summary

At enclosure opening the sky is clear with no wind, and 34% humidity
At the beginning of sxcientific operations the sky is clear, whit 1.0 seeing
Stuck of the readout in the MODS red channel. The SA sees the image in the mods computer, but it is not written in /newdata. Aferter one out of efforts spent in calling peaople and leaving them messages, we ask to switch instrument.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Elias_Rosa 11FE dual grating 1h Data acquired, but still don't saved at the end of the night. LOST the day after
Gilli4_A12 J1641 r-Sloan, i-sloan 1.5h, 1.5h partly within seen spec
Standard Sa112 r-Sloan, i-sloan, z-Sloan   airmas 1.18
Gilli4_A14 J2229 r-Sloan, i-sloan 30m, 30m within seeing specs, the last 4 frames with increasing backgorund

Jun 22-23

Summary

At the enclosure clear sky, no wind and 44.5% humidity
The Sa confirms we lost 1h of Elias-Rosa_47_DDT integrated yesterday night. The issue could come from the PINAME string that was too long and the computer was not able to write the corresponding header on the fits file.

Data Summary

Project_RUN Objects Filter ExpTime Notes
EliasRosa_47_DDT SN2011fe     COMPLETED
Israel_28_A4 CXOUJ204734 Dual 1h 30min Target completed
Israel_28_A5 CXOUJ215447 Dual 1h min Last 30min acquired during twilight
Topic revision: r1 - 23 Jun 2014, MarcoFumana
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback